OpenStudy (anonymous):

simplify radical 72??

OpenStudy (anonymous):

\[\Large \sqrt {36*2}=\sqrt{36}\sqrt{2}\] i hope now you can d this.

OpenStudy (anonymous):

do*

OpenStudy (anonymous):

6 radical 2??

OpenStudy (anonymous):

yup.